# Explainable Metrics for the Assessment of Neurodegenerative Diseases through Handwriting Analysis

### v1

@@ Abstract
Motor changes are early signs of neurodegenerative diseases (NDs) such as Parkinson’s disease (PD) and Alzheimer’s disease (AD), but are often difficult to detect, especially in the early stages. In this work, we examine the behavior of a wide array of explainable metrics extracted from the handwriting signals of 113 subjects performing multiple tasks on a digital tablet. The aim is to measure their effectiveness in characterizing and assessing multiple NDs, including AD and PD. To this end, task-agnostic and task-specific metrics are extracted from 14 distinct tasks. Subsequently, through statistical analysis and a series of classification experiments, we investigate which metrics provide greater discriminative power between NDs and healthy controls and among different NDs. Preliminary results indicate that the various tasks at hand can all be effectively leveraged to distinguish between the considered set of NDs, specifically by measuring the stability, the speed of writing, the time spent not writing, and the pressure variations between groups from our handcrafted explainable metrics, which shows $p$-values lower than 0.0001 for multiple tasks. Using various classification algorithms on the computed metrics, we obtain up to 87% accuracy to discriminate AD and healthy controls (CTL), and up to 69% for PD vs CTL.
1 Introduction
@@
Neurodegenerative diseases (NDs), such as Parkinson’s disease (PD) and Alzheimer’s disease (AD), have detrimental effects on specific regions of the brain, resulting in a gradual deterioration of cognitive, functional, and behavioral capabilities. This deterioration can be monitored, which is why handwriting, with its complex nature that involves cognitive, kinetic, and perceptual motor components, plays a significant role in the evaluation of NDs [1, 2]. Handwriting involves crucial brain regions related to learning and execution functions, including the cerebral cortex, basal ganglia, and cerebellum [3], which is why alterations in its patterns can serve as source of prominent biomarkers, indicative of both the presence and severity of the disease [4].
@@
Paper and digital handwriting tests [4, 5, 1] have been used to assess the presence of NDs through various signs. Previous works have been mainly focusing on handwriting for PD assessment [6, 7, 8, 9], with a handful starting to tackle AD assessment as well [10]. Subjects with PD often suffer from micrographia [11, 12], tremor [12], bradykinesia [12], and hypometria [13], a collection of signs collectively referred to as PD dysgraphia[1].
@@
The handwriting tests used for the possible assessment of NDs can be broken down into several categories that are used to collect various metrics. The first category is simple drawing tasks, including spirals and straight lines. Simple drawings can be used to assess trajectory, tremor, dimensionality, velocity, and acceleration [14]. The second category is simple writing tasks, including simple words and short sentences. When performed on a digital tablet, these tasks allow for the analysis of in-air movements and motor planning between words [14]. Collecting in-air movements is important because the time in-air results from a person planning out their next movement, which reflects their cognitive ability [5]. Additionally, in-air and on-surface movements were found to be statistically independent and can provide non-redundant information about the writer [15]. The in-air time is also indicative of hesitation and motor planning activities. Fluid handwriting, or handwriting with minimal in-air time, demonstrates the presence of an anticipated motor plan. At the same time, longer hesitations between writing might show that the writer needs to rethink their motor plan. The third category is complex tasks, which include complex drawing or paragraph copying. Complex drawing tasks and freewriting, as in the clock drawing task [16], allow for the detection of visual-spatial deficits by analyzing a person’s spatial organization skills. Copying tasks provide comparisons of writing variations with respect to different stimuli [17]. The use of a variety of these tasks provides a full profile of a person’s handwriting capabilities.
@@
In this study, we collected a new data set from 103 participants with various NDs, such as AD, Mild Cognitive Impairment (MCI) due to AD, PD, and various Parkinsonisms that are difficult to differentiate from PD and that we will name Parkinson’s Disease Mimics (PDM). This dataset comprises recordings of handwriting on a digital tablet across 14 specific tasks designed to assess neurodegenerative conditions. We extract explainable and quantitative endpoints, also called metrics, from 14 handwriting tasks and pinpoint key metrics capable of distinguishing between healthy controls and various NDs groups, including an active control group to ocmpare to PD participants: the PDM. We first explore a more general approach to obtain task-agnostic metrics, independently of their original purpose or content.
@@
Secondly, a task-dependant approach is adopted, to fully explore which traits of the participants can be used for discrimination, following the expected behavior that each group of participant is supposed to exhibit for each group of tasks.
@@
Furthermore, we use statistical analysis and machine learning classifiers to compare the prediction power of the extracted metrics, using binary comparisons between groups. The results obtained from that analysis highlight the possible value of these endpoints as digital biomarkers.
@@
  • The inclusion of several ND groups, including an active control group to compare to PD (the PDM group), where all precedent studies were limited to studying one or two NDs.
  • The use of Handwriting for AD and MCI assessment, usually reserved to PD assessment.
  • A large set of task-agnostic and task-specific metrics that show significant differences between groups of participants, following expected behaviors.
@@
The following sections of the manuscripts are organized as follows:
  • Section 2 presents the data collection pipeline and the main pre-processing stages.
@@
  • Section 3 introduces the task-agnostic metrics extracted from all tasks and the task-specific metrics designed for specific groups of tasks. Then, various methods used to gauge the significance of our metrics in the various tasks analyzed are described. Finally, we introduce the methods used for training classifiers for NDs assessment from the presented explainble metrics.
  • Section 4 shows and discusses our statistical and classification results.
  • Section 5 contains the main findings and conclusions of the study.
2 Materials
@@
In this section, we outline the data collection and preprocessing methods, along with the tasks evaluated in this research.
2.1 Data Collection
@@
The authors of this study collected a dataset from 113 participants [18, 19, 20, 21, 22, 23]. These participants were either categorized as healthy (referred to as CTL) or diagnosed and treated for a neurodegenerative disease by a specialist neurologist or geriatrician from Johns Hopkins Hospital. Participants diagnosed with NDs were categorized into one of several groups based on their specific diagnosis:
  • AD participants clinically diagnosed and treated by an AD expert geriatrician.
  • MCI participants presenting signs of MCI, evaluated by the same geriatrician, but without any AD diagnosis yet.
  • PD participants with clinically established diagnosis and being treated by a movement disorders neurologist.
@@
  • PDM participants presenting symptoms similar to those of PD, evaluated by the same neurologist, but with a diagnosis different from PD. This group includes patients initially diagnosed with Dementia with Lew Bodies, Fahr’s Disease, Ataxia, Corticobasal Syndrome, Corticobasal Degeneration, Cervical Dystonia, Gerstmann Schenker Schyuler Syndrome, Wilson Disease, MSA-Parkinsonian, and Essential Tremor. In this study, this group is considered an active control group against the PD group.
@@
All participants signed informed consent, and the data collection was approved by the Johns Hopkins Medical Institutional Review Board. Participants with PD continued their usual pharmacological treatment and took dopaminergic medication before the recording session. Handwriting samples were collected using a Wacom One 13 tablet with the associated pen. This setup enabled us to record the pressure applied by the pen on the tablet surface, as well as its position both on the tablet and in the air when in close proximity. Additionally, the recordings encompass the synchronized speech of participants and their ocular eye movements, although these two modalities were not utilized in this study. Participants completed 14 written tasks per session, detailed in Section 2.2, guided by a research assistant. We collected recordings for each participant for at least one session, while some participants were recorded up to three times, when available and interested, with each session spaced six months appart. The distribution of participants, with their age, gender, and total number of tasks recorded, is presented in Table 1 Participants classified as AD and MCI were combined into a single group, named AD in this study. Only a subset of the recorded participants were kept in that study, to balance as much as possible the age and gender of the participants. However, gender balancing for the AD group was not possible, given the low number of female participants, so we choose to keep the whole group.
Table 1: Distribution of the participants recorded in the dataset.
@@
2.2 Handwriting Tasks
In each session, we collected data from 14 handwriting tasks per participant. Due to various reasons, technical or human, the recordings of some tasks can fail, leading to an different number of participants for each task. Participants having one or more missing tasks are still included in the study. Within this section, we provide a comprehensive description of tasks, their origin (what each of those tasks were used for, originally), and elucidate the experimental hypotheses we formulated for each. Tasks such as Point and Spiral tasks involved participants performing stability exercises three times: alternating hands, then with their dominant hand while concurrently speaking. Writing and Drawing tasks prompted participants to replicate an item from the screen, from memory, and generate new content. Participants were asked to not touch the tablet with their arm or hand while writing. Figure 1 illustrates a sample from each task type, encompassing participants from the CTL, AD, PD, and PDM groups.
@@ 2.2.1 Point Tasks
2.2.2 Spiral Tasks
The spiral tasks are also designed to assess the motor capacities of the participants. Multiple studies require participants with PD to perform regular circular movement on paper or on tablet to assess their regularity or the presence of an eventual tremor [24, 25, 26]. The spiral test has also proven useful for AD assessment using a tablet [27]. The participants were asked to draw a spiral starting from the center of the tablet. More specifically, in the Spiral Dominant and Spiral Non-Dominant tasks, participants were asked to use their dominant and Non-Dominant hand respectively, and in the Spiral PaTaKa, they were asked to use their dominant hand, in addition, to perform a diadochokinetic task, i.e., repeat as fast as possible the syllables pa, ta, and ka in this exact order. However, our current focus lies solely on the handwriting data. We hypothesize that the simultaneous performance of tasks will make it more difficult for participants to use strategies to compensate for the motor impairments caused by NDs.
2.2.3 Writing Tasks
@@
Participants first engaged in the Copy Text task, wherein they were tasked with replicating a paragraph of text displayed on the screen. We then focus on the impact of multitasking by introducing the Copy Read Text task, where participants copied another text while simultaneously reading it out loud. Subsequently, participants’ spontaneous writing abilities were captured through the Freewrite task, prompting them to write freely one or two sentences that must not contain any personal information. Lastly, a different cognitive function was explored via the Numbers task, wherein participants were tasked with solving eight simple arithmetic operations involving the sum of two natural numbers not exceeding a total of 20, while saying the results of the sums out loud. It is worth noting that the Numbers task may be analyzed differently from the writing tasks due to its repetitive nature, necessitating a distinct analytical approach.
2.2.4 Drawing Tasks
The Drawing tasks are also designed to assess the cognitive and motor capacities of the participants.
First, in the Copy Image task, we asked participants to copy an image displayed on the screen. The image is shown in the figure 2
@@ 2.2.4 Drawing Tasks
We further examined the impact of memory on this behavior by introducing the Copy Image Memory task. After copying the image in the Copy Image task, participants were then asked to reproduce it from memory approximately five minutes later.
Subsequently, participants were asked to perform two tasks: First, the Draw Clock task, which looked at their proficiency in drawing all the required elements of a clock at eleven past ten, used for AD assessment in the MoCA test [29] and for PD assessment [16], and for discrimination between neuro-degenerative diseases [30].
@@ 2.3 Data Curation
3 Methods
This section describes the metrics we computed for each task and the approach used to assess their significance between the various experimental groups considered. We meticulously crafted a set of interpretable metrics derived from the raw handwriting signals collected during the various tasks of interest. Our analysis of the various tasks is twofold: firstly, we scrutinize the dynamic properties of handwriting, which are common to all tasks. Secondly, we evaluate their performance in specific tasks, such as stability for point tasks and proximity between drawing and target for copying tasks. Dynamic analysis is conducted across all relevant tasks to extract overarching metrics concerning handwriting, which will be used to measure certain behavioral aspects of each disease. Conversely, task-specific metrics provide deeper insights into particular aspects of the disease by investigation of a limited set of tasks at a time.
3.1 Metrics computation pipeline
@@
We computed various metrics from the clean files, targeting various cognitive and behavioral elements. The metrics can be categorized into two distinct sets:
@@
    Task-agnostic (see Section 3.2): these metrics measure primarily dynamic properties of the signal such as speed, noise, and basic ratios.
@@
    Task-specific (see Section 3.3): these metrics are designed for specific tasks or groups of tasks.
3.2 Task-Agnostic Metrics
Task-agnostic metrics quantify the dynamic aspects of the drawings. They are computed using various signal-processing techniques. The extraction process, described in Figure 3, consists of three distinct steps:

### v2

@@ Abstract
Motor dysfunction is a common sign of neurodegenerative diseases (NDs) such as Parkinson’s disease (PD) and Alzheimer’s disease (AD), but may be difficult to detect, especially in the early stages. In this work, we examine the behavior of a wide array of explainable metrics extracted from the handwriting signals of 113 subjects performing multiple tasks on a digital tablet, as part of the Neurological Signals dataset. The aim is to measure their effectiveness in characterizing NDs, including AD and PD. To this end, task-agnostic and task-specific metrics are extracted from 14 distinct tasks. Subsequently, through statistical analysis and a series of classification experiments, we investigate which metrics provide greater discriminative power between NDs and healthy controls and amongst different NDs. Preliminary results indicate that the tasks at hand can all be effectively leveraged to distinguish between the considered set of NDs, specifically by measuring the stability, the speed of writing, the time spent not writing, and the pressure variations between groups from our handcrafted explainable metrics, which shows $p$-values lower than 0.0001 for multiple tasks. Using various binary classification algorithms on the computed metrics, we obtain up to 87% accuracy for the discrimination between AD and healthy controls (CTL), and up to 69% for the discrimination between PD and CTL.
Index Terms:
Handwriting Analysis, Alzheimer’s Disease, Parkinson’s Disease, Digital Biomarkers, Neurological Signals
I Introduction
@@
Neurodegenerative diseases (NDs), such as Parkinson’s disease (PD) and Alzheimer’s disease (AD), have detrimental effects on shared and disparate regions of the brain, resulting in a gradual deterioration of cognitive, motor, and behavioral function. Numerous attempts have been made to track dysfunction utilizing non-invasive, wearable or digital sensors. Handwriting, in particular, with its complex syngergism between cognitive, kinetic, and perceptual motor components, is therefore an obvious avenue to evaluate NDs [1, 2]. Successful handwriting utilizes cortical and subcortical brain regions pertaining to learning, cognitive control, praxis, vigilance and coordination across the cerebral cortex, basal ganglia, and cerebellum [3]. Alterations in its patterns can serve as a source of prominent digital markers, indicative of both the presence and severity of the disease [4], as exemplified by micrographia being a common early sign of PD [5].
@@
Paper and digital handwriting tests [4, 5, 1] have been used to assess the presence of NDs through the dynamic and spacial characteristics of handwriting across sets of tasks. The bulk of prior work studied PD as micrographia was an obvious target [6, 7, 8, 9]. Subjects with PD often suffer from micrographia [10, 11], tremor [11], and bradykinesia [11], a collection of signs collectively referred to as PD dysgraphia[1]. More recently, numerous reports have begun to tackle AD detection through handwriting test, as the condition has grown more common [12].
@@
The handwriting tests used for assessing NDs can be categorized into several subtypes, dependent upon complexity and specific task instructions. The first category consists of simple drawing shapes, such as spirals and straight lines. These simple drawings readily evaluate trajectory, tremor, dimensionality, velocity, or acceleration [13]. The second category involves writing words or short sentences. When conducted on a digital tablet, these tasks permit analysis of in-air movements and motor planning between words [13]. The collection of in-air movements is crucial, as the time spent in-air reflects a person’s preparation and planning for subsequent movements, implicitly indicative of their cognitive processesing [5]. Additionally, in-air and on-surface movements are statistically independent and provide distinct information about the writer [14]. A prolonged in-air time may be indicative of hesitations or impaired motor planning and sequencing. In contrast, fluid handwriting execution with minimal in-air time, demonstrates the presence of an established motor plan. The third category encompasses complex tasks, including drawing complex figures or transcribing paragraphs. The clock drawing task is an established modality [15] to detect cortical and subcortical dysfunction across several NDs that reflect organization and visuospatial skills; or visuoconstruction deficits [16]. Copying tasks provide comparisons of writing variations with respect to different stimuli [17]. The use of a variety of these tasks provides a full profile of a person’s handwriting capabilities.
@@
In this study, we collected a new data set from 103 participants with several NDs: AD, Mild Cognitive Impairment (MCI), PD and Related Disorders (the latter term inclusive of atypical PD and secondary Parkinsonism in which symptoms were initially ascribed to PD but later concluded to be an alternative condition, which we will term [PDM]). This dataset comprises recordings of handwriting on a digital tablet across 14 specific tasks specifically designed to assess all aforementioned NDs. We extract explainable and quantitative endpoints, also called metrics, from 14 handwriting tasks and pinpoint key metrics that distinguish between healthy controls and the NDs subgroups, including an active control group to compare against PD (i.e., PDM). We first explore a more general approach to obtain task-agnostic metrics, independently of their original purpose or content.
@@
Secondly, a task-dependent approach is adopted to fully explore which traits of the participants can be used for discrimination, following the expected behavior that each group of participants is supposed to exhibit for each group of tasks.
@@
Furthermore, we use statistical analysis and machine learning classifiers to compare the prediction power of the extracted metrics using binary comparisons between groups. The results obtained from that analysis highlight the possible value of these endpoints as digital biomarkers.
@@
  • The inclusion of several ND groups, including an active control group to compare to PD (i.e, PDM group), where all precedent studies were more limited in scope.
  • The use of Handwriting for AD and MCI assessment was previously primarily reserved for PD assessment.
  • A large set of task-agnostic and task-specific metrics that show significant differences between groups of participants following expected behaviors.
@@
The sections of the manuscripts are organized as follows: Section II presents the data collection pipeline and the main pre-processing stages. Section III introduces the task-agnostic metrics extracted from all tasks and the task-specific metrics designed for specific groups of tasks. Then, various methods used to gauge the significance of our metrics in the various tasks analyzed are described. Finally, we introduce the methods used for training classifiers for NDs assessment from the presented explainable metrics. Section IV shows and discusses our statistical and classification results. Section V contains the main findings, discussions, limits and conclusions of the study.
II Materials
@@
In this section, we outline the data collection and preprocessing methods and the tasks evaluated in this research.
II-A Data Collection
@@
The authors of this study collected a dataset from 113 participants from a larger digital biomarker cohort, NeuroLogical Signals, which our group has previously reported [18, 19, 20, 21, 22, 23]. These participants were either categorized as healthy (CTL) or diagnosed and treated for a ND by a subspecialist neurologist or geriatric psychiatrist from Johns Hopkins Hospital. Participants diagnosed with NDs were categorized into aforementioned groups based on disease specific clinical diagnostic criteria diagnosis:
  • AD participants were enrolled from a University AD Center of Excellence, treated by an expert geriatric psychiatrist with an AD diagnosis based upon 2011 criteria [24]. However, most of these subjects were enrolled in serum or CSF biomarker studies confirming their diagnosis reflecting later revisions to diagnostic criteria [25].
  • MCI participants were characterized by established criteria, not yet meeting multidomain impairment or disability consistent with AD [26].
  • PD participants met international recognized diagnostic criteria at a ”clinically established” [27] level of certainty and were followed by a movement disorders neurologist.
@@
  • PDM participants presented with one or more core symptoms of PD, meeting a ”Probable PD” diagnosis but which was later revised based on evolving characteristics or supportive testing. This is a mixed, active comparator group includes patients later diagnosed with Dementia with Lewy Bodies, Fahr’s Disease, Ataxia, Corticobasal Syndrome, Corticobasal Degeneration, Cervical Dystonia, Gerstmann Schenker Schyuler Syndrome, Wilson Disease, MSA-Parkinsonian, and Essential Tremor. In this study, this group is considered an active control group against the PD group.
@@
All participants signed informed consent, and the data collection was approved by the Johns Hopkins Medicine Institutional Review Board. Participants with PD continued their usual pharmacological treatment and took dopaminergic medication before the recording session. Handwriting samples were collected using a Wacom One 13 tablet [^1] https://estore.wacom.com/en-us/wacom-one-13-touch-dth134w0a.html with the associated pen. This setup enabled us to record the pressure applied by the pen on the tablet surface and its position both on the tablet and in the air when in close proximity. Note the recording session also included collection of saccadometric and synchronized speech, although these two modalities were not utilized in this study. Participants completed 14 written tasks per session, detailed in Section II-B, guided by a research assistant. We collected recordings for each participant for at least one session, while 21 participants were recorded up to three times, when available and interested, with each session spaced six months apart. The handwriting data collected is part of a larger, multimodal dataset that is yet to be published. The distribution of participants, with their age, gender, and total number of tasks recorded, is presented in Table I Participants classified as AD and MCI were combined into a single group, named AD in this study. Only a subset of the recorded participants were kept in that study, to balance as much as possible the age and gender of the participants. However, given the low number of female participants, gender balancing for the AD group was not possible, so we chose to keep the whole group.
TABLE I: Distribution of the participants recorded in the dataset.
@@
II-B Handwriting Tasks
In each session, we collected data from 14 handwriting tasks per participant. Due to myriad reasons, technical or human, the recordings of some tasks can fail, leading to a variable number of participants for each task. Participants having one or more missing tasks are still included in the study. Within this section, we provide a comprehensive description of tasks, their origin (what each of those tasks were used for, initially), and outline the experimental hypotheses we formulated for each. Tasks such as Point and Spiral tasks involved participants performing stability exercises three times: alternating hands, then with their dominant hand while concurrently speaking. Writing and Drawing tasks prompted participants to replicate an item from the screen, from memory, and generate new content. Participants were asked not to touch the tablet with their arm or hand while writing. Figure 1 illustrates a sample from each task type, encompassing participants from the CTL, AD, PD, and PDM groups.
@@ II-B1 Point Tasks
II-B2 Spiral Tasks
The spiral tasks are also designed to assess the dynamic motor control and coordination of the participants. Several clinical rating scales [28, 29] require participants with PD to perform regular circular movement on paper or on tablet to assess their regularity or the presence of an eventual tremor [30, 31, 32]. The spiral test has also proven useful for AD assessment using a tablet [33]. The participants were asked to draw a spiral starting from the center of the tablet. More specifically, in the Spiral Dominant and Spiral Non-Dominant tasks, participants were asked to use their dominant and Non-Dominant hand respectively, and in the dual-task Spiral PaTaKa, they were asked to use their dominant hand, and simultaneously, to perform a spoken diadochokinetic task, i.e., repeat as fast as possible the syllables pa, ta, and ka in this exact order. However, our current focus lies solely on the handwriting data. We hypothesize that the simultaneous performance of tasks will make it more difficult for participants to use strategies to compensate for the motor impairments caused by NDs.
II-B3 Writing Tasks
@@
Participants first engaged in the Copy Text task, wherein they were tasked with replicating a paragraph of text displayed on the screen. We then focus on the impact of multitasking by introducing the Copy Read Text task, where participants copied another text while simultaneously reading it aloud. Subsequently, participants’ spontaneous writing abilities were captured through the Freewrite task, prompting them to write freely one or two sentences that must not contain any personal information. Lastly, we exploring handwriting a different context, the Numbers task, wherein participants were tasked with solving eight simple arithmetic operations summing two natural numbers not exceeding a total of 20, while saying the results of the sums out loud. It is worth noting that the Numbers task may be analyzed differently from the writing tasks due to its repetitive nature, necessitating a distinct analytical approach.
II-B4 Drawing Tasks
The Drawing tasks are also designed to assess the cognitive and visuo-constructive capacities of the participants.
First, in the Copy Image task, we asked participants to copy an image displayed on the screen. The image, is a complex figure in the style of those used in the Complex Figure Task [35]: asymmetric, non-patterned with internal and external details, as shown in the figure 2
@@ II-B4 Drawing Tasks
We further examined the impact of memory on this behavior by introducing the Copy Image Memory task. After copying the image in the Copy Image task, participants were then asked to reproduce it from memory approximately five minutes later, thus assessing visual memory and construction abilities.
Subsequently, participants were asked to perform two tasks: First, the Draw Clock task, which looked at their ability to draw all the required elements of a clock at an instructed time [36]. The Clock Draw Task is another established visuoconstructive screen studied in PD [16] and AD [37] which has been utilized to distinguish ND [38] and may be interpreted in a number of ways.
@@ II-C Data Curation
III Methods
This section outlines the metrics calculated for each task and the methods used to assess their significance among different experimental groups. We developed interpretable metrics from raw handwriting signals collected during various tasks. Our analysis has two aspects: first, we examine dynamic handwriting properties common to all tasks; second, we evaluate performance in specific tasks, such as stability in point tasks and proximity between drawing and target in copying tasks.
III-A Metrics computation pipeline
@@
We computed various metrics from the clean files, targeting prioritizing motor, cognitive or behavioral elements. The metrics can be categorized relative to or irrespective of the task:
@@
    Task-agnostic (see Section III-B): these metrics measure primarily dynamic properties of the signal such as speed, noise, and basic ratios.
@@
    Task-specific (see Section III-C): these metrics are designed for specific tasks or groups of tasks.
III-B Task-Agnostic Metrics
Task-agnostic metrics quantify the dynamic aspects of the drawings. They are computed using several signal-processing techniques. The extraction process, illustrated in Figure 3, may be parsed: